OpenStudy (anonymous):

What is the slope of the line 5x + 2y = 10?

OpenStudy (anonymous):

you need to get the equation in the proper form. First you need to subtract 5x from both sides of the equation giving you 2y= -5x +10. Then you will need to divide 2 from both sides of the equation so that y is by itself. y= -5/2 x +10/2. Then simplify. y = -5/2 x + 5. Your slope will be -5/2. This is true because y = mx + b, where m is the slope of your line.

OpenStudy (raden):

alternative : if given a straight line equation : ax + by = c then the slope always m = -a/b